# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

### v2

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes. Evaluating both vision-language models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available. Our best-performing transformer-based fine-tuned model reaches 83% f-score, outperforming gpt4 by 23%. Also, we demonstrate how this model can be used to uncover trends during wildfires. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@
Braeden Sherritt Carleton University                        Isar Nejadgholi National Research Council Canada                        Marzieh Amini Carleton University
@@ 1 Introduction
To foster research in this area, Alam et al. (2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. The HumAID dataset by Alam et al. (2021) also contains wildfire data from Canada, but only a couple of thousand samples from 2016.
Therefore, wildfires in Canada remain underexplored in social media analysis research. To address this limitation, we focus our work on gathering social media data specifically related to Canadian wildfires, as they are becoming increasingly frequent and severe due to the adverse effects of climate change.
@@
Upon initial inspection of our Canadian wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of this data. To address this, we devised a novel taxonomy of twelve categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts. Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
@@
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several VLMs, as well as the feasibility of training custom classifiers. Our best-performing classifier is a fine-tuned model, which uses RoBERTa Liu (2019) and Vision Transformer (ViT) models Ofli et al. (2020); Basit et al. (2023) to jointly process text and image parts of tweets. While WildFireCan-MMD is sourced from data posted in 2022-2024, we collected an additional set of unlabeled tweets across seven years (2018-2025). We then used our best-performing classifier to analyze this data and uncover trends and patterns of social media use in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at WildFireCanMMD-Github [^1] Please contact the authors for more information about data sharability., under CC-BY-NC license. The contributions of this study are summarized as follows:
@@
  • We devised a twelve-category taxonomy through multimodal topic modeling of Canadian wildfire-related user-generated content, and introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, manually labeled based on our taxonomy.
@@
  • We compared custom multimodal transformer models trained on WildFireCan-MMD with several VLMs and classical methods. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
@@
  • We collected an additional 46,279 posts from X, spanning across seven years, used our trained classifier to label them, and reported the uncovered trends and patterns of social media use during wildfires in Canada.
@@ 2.1 Social media analysis for disaster response
Our work is motivated by this literature and focuses on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
@@ 2.2 Multimodal classification for disaster response
Alam et al. (2018) addressed the shortage of multimodal labeled social media data for disaster response by introducing CrisisMMD, a widely used dataset containing annotated tweets and images from various natural disasters. Later extended by datasets like HumAID Alam et al. (2021), CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Basit et al. (2023); Koshy and Elango (2023); Shetty et al. (2024).
While previous works such as CrisisMMD used a generalized scheme of annotations across various types of different disasters, our work develops a wildfire-specific taxonomy of twelve categories and introduces WildFireCan-MMD, a new Canadian wildfire dataset.
@@
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT Lan et al. (2020) as a text embedder, and DeIT Touvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018). We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
@@ 3.1 Data Collection
To construct our dataset, we utilized X’s Pro-Tier paid API to collect a diverse set of posts surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from tweets posted between April and June of each respective year. Data collection was conducted using a hashtag-based search strategy. Table A.1 presents the full queries used for data collection. To curate the hashtags, we started with seed hashtags and used an iterative process of adding new hashtags by collecting sample tweets. This approach was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, we identified and utilized hashtags that, through iterative experimentation, were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated.
@@ 3.2 Taxonomy Development
To examine the content of the tweets and discover the hidden latent semantics of our data, we used BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). Further explanation of the BERTopic clustering method are provided in Appendix B.
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting number of topics to sort the data into. When the number of clusters were unspecified, BERTopic had created over 100 fine-grained topics. This was relatively large to manage, and difficult to determine a humanitarian category for each fine-grained topic. We gradually decreased the number of topics, and for each number, we manually inspected the representative visual samples of topics and keywords generated by BERTopic. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about infrastructure, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, including representative images and corresponding keywords, generated by BERTopic, can be seen in Figure B.1. We then manually inspected the topics and selected 12 classes that were informative and consistently appeared across a varying maximum number of clusters to be considered designated categories for this dataset.
The full list of categories and their descriptions is given in Table 1. We also include an “Other” to accommodate the examples that do not fit into any of these 12 categories. The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
@@
Evacuees Information relating to evacuees, their movements, needs, location. 242
Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1070
General Information General info about the wildfire situation, such as total hectares burned. 186
Preparedness Information for the general public to prepare themselves and their property. 265
Weather Reports Information relating to the weather with a specific location mentioned. 316
Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 620
Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 378
Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 293
Political Posts directed towards political figures or parties (excluding situation updates). 322
Support Mental health and financial support, temporary housing for livestock. 164
Insurance Information relating to insurance, employment insurance, and EI benefits. 142
Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 116
Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 569
@@ 3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Over the course of labeling, we ensured that the data did not contain personal information. Labeling was carried out in-house by one of the authors using Label Studio [^2] https://labelstud.io/. To ensure the reliability of the labeling process, we first discussed the definitions and boundaries of each category among the research team and reached a shared understanding. Through this discussion, we determined that the task was not largely subjective, as the categories are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image). Therefore, we did not consider multi-annotator labeling necessary. Examples of labeled tweets are presented in Figure 1. The size of labeled classes are presented in Table 1.

### v3

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. In this work, we focus on multimodal wildfire social media data, which, although existing in current datasets, is currently underrepresented in Canadian contexts. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes. We evaluate zero-shot vision-language models on this dataset and compare their results with those of custom-trained and baseline classifiers. We show that while baseline methods and zero-shot prompting offer quick deployment, custom-trained models outperform them when labelled data is available. Our best-performing custom model reaches 84.48±0.69% f-score, outperforming VLMs and baseline classifiers. We also demonstrate how this model can be used to uncover trends during wildfires, through the collection and analysis of a large unlabeled dataset. Our dataset facilitates future research in wildfire response, and our findings highlight the importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@
Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,MarziehAmini@cunet.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca
@@ 1 Introduction
To foster research in this area, Alam et al. (2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising 1589 samples, and is specific to the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset Alam et al. (2021) are limited to a single month in 2016, as well as another month of California samples from 2018. Therefore, wildfires in Canada remain underexplored in social media analysis research.
@@
To address this limitation, we focus our work on gathering social media data specifically related to Canadian wildfires, as they are becoming increasingly frequent and severe due to the adverse effects of climate change. Upon initial inspection of our Canadian wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of this data. To address this, we devised a novel taxonomy of twelve categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts. Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
@@
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several Vision-Language Models (VLMs), custom deep-learning classifiers, as well as classical approaches, as a minimum achievable performance. Our best-performing classifier utilizes RoBERTa Liu (2019) and Vision Transformer (ViT) models to jointly process the text and image components of tweets. A custom transformer-based fusion module is designed to enable the modalities to interact deeply. While WildFireCan-MMD is sourced from data posted between 2022 and 2024, we collected an additional set of unlabeled tweets spanning seven years (2018-2024). Using our best-performing classifier, we analyze this unlabeled data and uncover trends and patterns of social media use during wildfires in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at [Anonymized GitHub], under CC-BY License. The contributions of this study are summarized as:
@@
  • We collected 4,688 wildfire-related posts from X in Canada from 2022 to 2024 and devised a twelve-category taxonomy using unsupervised topic modeling and human analysis. We introduce WildFireCan-MMD, the first Canada-specific multimodal wildfire social-media dataset, annotated by three annotators.
@@
  • We propose a custom multimodal transformer model, trained on WildFireCan-MMD and compare it with several VLMs and classical methods. The fine-tuned custom classifier outperforms zero-shot VLMs significantly; our model achieves an f-Score of 84.48±0.69%.
@@
  • Additionally, we collected 46,279 posts from X from 2018 to 2024. We used our trained classifier to label them and reported the trends and patterns of social media use during wildfires in Canada. Results show that retrospective analysis using a classifier can uncover trends which relate to real situational updates.
@@ 2.1 Social media analysis for disaster response
Our work is motivated by this literature as well as the literature shown in Appendix G which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
@@ 2.2 Multimodal classification for disaster response
The study by Alam et al. (2018) released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes). CrisisMMD contains thousands of tweets labeled for categories like “infrastructure and utility damage” or “rescue volunteering or donation effort”.
Additionally, Alam et al. (2018) later released the HumAID Alam et al. (2021) dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019). It is similarly annotated for emergency response categories (e.g. casualties, infrastructure damage).
@@
There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex Olteanu et al. (2015) and CrisisMMD Alam et al. (2018) datasets Wu et al. (2022). CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Ofli et al. (2020); Khattar and Quadri (2022); Basit et al. (2023); Koshy and Elango (2023); Rezk et al. (2023) and many others. Therefore, there is a notable gap in the in the availability of multimodal wildfire datasets.
In the context of Canada, where the most common and destructive natural disaster is wildfires, CrisisMMD is lacking in data. Although it contains  18k samples from other disasters such as hurricanes, earthquakes and floods, the amount of wildfire data is proportionally small (1589 samples). Additionally, this data is only from the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016, as well as another month of California samples from 2018. Therefore, wildfires in Canada remain underexplored in social media analysis research.
To address this limitation, this work focuses on gathering social media data related to Canadian wildfires, and proposes WildFireCan-MMD, a new Canadian wildfire social-media dataset with a wildfire-specific taxonomy 12 categories.
@@
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT Lan et al. (2020) as a text embedder, and DeIT Touvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018).
Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix G. Most previous studies use a CNN for the image modality and then a BERT-based transformer for the text modality. Other works sometimes use ViT for the image modality. Concatenative fusion of the two modalities into a series of linear layers seems to be the most common approach, with other works using multiplicative or additive fusion. Cross-attention fusion of the two modalities seems to be another popular approach that seems to give promising results. The approach given by Pranesh (2022) proposes a transformer-like structure to fuse the two modalities. We build on the this previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
@@ 3.1 Data Collection
To construct the dataset, X’s Pro-Tier paid API was utilized to collect a diverse set of posts, including images and text, surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from the wildfire season (April-September) of each respective year. Data collection was conducted using a hashtag-based search strategy as seen in other works Alam et al. (2018); Gupta and Hewett (2020); Cruickshank and Carley (2020). Appendix A presents the full queries used for data collection. To curate the hashtags, basic hashtags such as "#wildfire, #forestfire" were first used. Then, through an iterative process of adding new hashtags by collecting sample tweets, more relevant hashtags were discovered. This approach of using hashtags was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, hashtags identified through iterative experimentation were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region from which the tweet originated.
@@ 3.2 Taxonomy Development
To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst (2022). Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic. The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15. Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords. Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics. Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased. For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected. Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear. Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders. Less useful posts were also noticed, such as advertisements and political debates/complaints. Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix C.
The topics were inspected and 12 classes were decided upon. The full list of classes and their descriptions is given in Table 1. To ensure the reliability of the labeling process, the definitions and boundaries of each category were discussed among the research team and a shared understanding was reached. Through this discussion, it was determined that the classes are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image). This discussion, along with unsupervised topic modeling (BERTopic) were used to find strong topics which were then later assigned to a class. Also, the selection of classes was guided by the topics created by BERTopic, observations made through manual annotation, and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster. An ‘Other’ class is added to accommodate the samples that do not fit into any of the 12 classes.
@@
Evacuees Information relating to evacuees, their movements, needs, location. 252
Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1128
General Information General info about the wildfire situation, such as total hectares burned. 170
Preparedness Information for the general public to prepare themselves and their property. 264
Weather Reports Information relating to the weather with a specific location mentioned. 296
Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 669
Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 356
Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 264
Political Posts directed towards political figures or parties (excluding situation updates). 329
Support Mental health and financial support, temporary housing for livestock. 178
Insurance Information relating to insurance, employment insurance, and EI benefits. 158
Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 117
Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 507
@@ 3.3 Data Labeling
Labeling was carried out in-house by three independent annotators using Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 1. Examples of labeled tweets are presented in Figure 1.
Each image-text pair was given a single label. In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample. For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’. In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone. If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided. The size of labeled classes is presented in Table 1.
Annotation was done by three independent annotators, all authors of this paper. The annotators are male, 20-30 year old, fluent in English and reside in Canada. First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes. Over the course of labeling, samples were flagged that contained personal information.
A vote between the labels from the three annotators decided the final labels. Annotation agreement scores can be seen in Table 2. A majority agreement was reached on 88.4% of the labels. Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels. From the results, annotators 1 & 2 had the best agreement scores. This is further demonstrated by Cohen’s and Fleiss’ Kappa metrics, which indicate that a moderate-to-good level of agreement is achieved between the independent annotators. In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators. Additional statistics can be found in Appendix B.
Table 2: Annotation agreement statistics.
Metric Agreement
Majority Agreement (2 same) 88.4%
Full Agreement (all same) 45.1%
Vote between all/annotator 1 78.7%
Vote between all/annotator 2 77.7%
Vote between all/annotator 3 65.8%
Cohen’s Kappa annotator 1-2 63.5%
Cohen’s Kappa annotator 1-3 50.6%
Cohen’s Kappa annotator 2-3 49.2%
Fleiss’ Kappa 54.3%